OpenStudy (anonymous):

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path.The bicyclist heading north is riding 4 km/hour faster than the bicyclist heading south. At 10:15, they are 40 km apart. Find the two bicyclists’ rates

OpenStudy (kropot72):

Let the speed of the south-bound bicyclist = s Then speed of north-bound bicyclist = s + 4 The separation speed = s + s + 4 = 4 + 2s ...............(1) Distance travelled in 1.25 hours = 40 km Therefore separation speed = 40/1.25 ..............(2) The two values of separation speed in (1) and (2) must be equal. 4 + 2s = 40/1.25 5 + 2.5s = 40 ..............(3) Now you need to solve equation (3) to find the value of s. Speed of north-bound bicyclist = s + 4

OpenStudy (anonymous):

need answer

OpenStudy (anonymous):

yall suck